OpenStudy (anonymous):

Helppppppp :( For the following function, h , determine the functions, f and g so that : h(x)= f(g(x)) If h(x)=(8x+8/11x-11)^7 then f(x) g(x)

OpenStudy (mertsj):

f(x)=x^7 g(x)=(8x+8)/(11x-11)

OpenStudy (anonymous):

so then do i plug in the equation in x^7?

OpenStudy (mertsj):

I thought the problem asked you what the functions f and g are

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

can you show me what the next steps are? i'm a bit confused..

OpenStudy (mertsj):

The problem told you that the f(g(x))=

OpenStudy (anonymous):

steps

OpenStudy (anonymous):

i know what the problem said but can you show me?

OpenStudy (mertsj):

\[(\frac{8x+8}{11x-11})^7\]

OpenStudy (anonymous):

okay.. i did that

OpenStudy (mertsj):

Then it asked you what the f(x) is and what the g(x) is. So can you see that the f(x)=x^7 and you replace x with (8x+8)/(11x-11) that you will get the f(g(x))? There is really no work to show. It's just to demonstrate your understanding of function composition.

OpenStudy (anonymous):

there is work to show

OpenStudy (anonymous):

how can they give us a problem without any work???

OpenStudy (mertsj):

Supposing they ask you what 4 x 3 is. Would there be work to show?

OpenStudy (anonymous):

after you found f(x)= x^7 when you plug in 8X+8etc etc.... what is the number or answer from there?

OpenStudy (mertsj):

I'm thinking you don't know what the problem is asking.

OpenStudy (anonymous):

of course I can. I can show 4 columns and 3 rows of apples to equal to 12!

OpenStudy (anonymous):

i do. I submitted the "answer" of not showing the work.... hint hint wrong

OpenStudy (anonymous):

there is work. its math buddy

OpenStudy (mertsj):

The problem says if h(x)= ... Then f(x)= and you should write x^7 g(x)= and you should write (8x+8)/(11x-11)

OpenStudy (anonymous):

yes i did

OpenStudy (mertsj):

And?

OpenStudy (mertsj):

@satellite73 Maybe satellite can help you with your work. He's really smart.

OpenStudy (anonymous):

I see...

OpenStudy (anonymous):

thanks so much :)

OpenStudy (mertsj):

yw. sorry

OpenStudy (anonymous):

lol no worries, again thank you.